32. /5x-1/less than and equal to 14
the lines mean absoloute value
Solve each absolute value inequality and graph it.

Answer by Nate(3500)

|5x - 1| <= 14
5x - 1 <= 14 and 5x - 1 >= -14
5x <= 15 and 5x >= -13
x <= 3 and x >= -13/5
<----(-13/5)====(3)---->